Preet is 5 years older than Ben. She is also 7 years older than Jessy. The sum of their age 21. How old is each person?

Answer by wangsol(2)

Preet is 5 years older than Ben:
P=B+5 ------eq. 1
She is also 7 years older than Jessy:
P=J+7
J=P-7 ------eq. 2
The sum of their age 21:
P+B+J=21 ------eq. 3
eq. 1 in eq. 2:
if P=B+5
then J=B+5-7 ------eq. 4
substitute eq. 1 & eq. 4 in eq. 3:(solve for B)
B+5+B+B+5-7=21
(B+B+B)+(5+5-7)=21
3B+3=21
3B=21-3
3B=18
B=18/3
B=6
substitute B=6 in eq. 1 & eq. 4:
P=6+5=11
J=6+5-7=4
therefore:
P=11, B=6, J=4